Suppose you bought 700 shares of stock at an initial price of $43 per share. The stock paid a dividend of $0.40 per share during the following year, and the share price at the end of the year was $44. a. What is the capital gains yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) b. What is the dividend yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) c. What is the total rate of return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

a.Capital gains yield=(End price-Beginning price)/Beginning price

=(44-43)/43

=2.33%(Approx).

b.Dividend yield=Dividend/Beginning price

=(0.4/43)

=0.93%(Approx).

c.Total rate of return=Capital gains yield+Dividend yield

=(2.33+0.93)

=3.26%(Approx).